# The Wisdom of a Crowd of Brains: A Universal Brain Encoder

### v2

@@ A.2 Training and Lora Modification
The Lora adaptation  (Hu et al., 2021) is done by adding learned low rank matrices to weights in the original network. In the original paper the value projection and query projection weights ($W_{v}$,$W_{q}$) in the self attention block are modified. We only modify the output projection weights($W_{o}$) of the self attention block.
@@ Appendix C Statistical Tests
Additionally, we tested the prediction correlations of individual voxels (across all subjects) against a null hypothesis. The null hypothesis was based on the distribution of correlations between pairs of independent Gaussian random vectors of size N = 982, matching the number of image-fMRI pairs in the test set. The one-sided statistical significance was estimated by comparing the predicted voxel correlations for each subject on the test set with those from the null distribution. We applied a statistical threshold of P $<$ 0.05, correcting for multiple comparisons using the FDR procedure. Our results indicate that the majority of voxel predictions by the Universal Encoder are statistically significant, with an average of 98% across all subjects. Detailed results in Table S1.
@@ Appendix D Additional Figures
Figure S8: Comparison against additional Encoder models. Pearson correlation between predicted and ground-truth fMRI of Subject1 from NSD dataset. The figure presents comparisons across three prominent encoders from different papers (Gaziv et al., 2022; Wang et al., 2023; Takagi & Nishimoto, 2023), as well as the universal encoder trained on both single-subject and multi-subject data. (a) Shows the median values along with the 25th and 75th percentiles across the five models. (b) Compares correlation scores across different coarse brain regions (ROIs) for the same five models. .
@@
Figure S11: Retrieval Results: Our universal encoder trained on multiple subjects significantly improves retrieval accuracy over both subject-specific models as evidenced by the calculated p-value. We conducted 10 retrieval experiments, wherein we randomly sampled 999 different distracted images for each “query” fMRI. Results are compared across three models: (i) ”Baseline Encoder” – the encoder of (Gaziv et al., 2022) trained separately on each subject, (ii) ”Universal Encoder - Single Subject” – our architecture trained separately on each subject, and (iii) ”Universal Encoder - Multiple Subjects” – our model trained on combined data from 8 subjects. In (a) we present the mean top-1 accuracy (across the 10 experiments) along with the standard deviation for each model and subject. (b) depicts the mean top-5 accuracy along with the standard deviation for each model and subject.
Figure S12: MSE Evaluation - The Wisdom of a Crowd of Brains. By aggregating data from multiple subjects, our Universal-Encoder improves the encoding of any individual subject. We present results calculating the MSE between predicted and ground-truth fMRI (median value with 25th and 75th percentiles). We compared 3 models: (i) The ”Baseline” single-subject encoder of (Gaziv et al., 2022), (ii) ”Universal Encoder - single subject” – our architecture trained on each subject separately, (iii) ”Universal Encoder - multiple subjects” – our model trained on data from 8 subjects.
@@
Figure S13: Qualitative Evaluation of Image Reconstruction from fMRI. Visual comparison of image reconstruction from fMRI for Subject 1 from the NSD dataset. The figure showcases the ground truth (GT) image seen by the subject, alongside reconstructions generated using the fMRI-to-Image decoding model (MindEye (Scotti et al., 2024a)) trained on Subject 1’s fMRI data. Both the decoder and our Universal Encoder were trained on the same dataset and evaluated on the same test set. Reconstructions were evaluated on the test set using two inputs: (1) ”Reconstruction from GT fMRI” where the decoder processes the real fMRI recordings of the test images, and (2) ”Reconstruction from our encoded fMRI”, where the test images are first encoded into predicted fMRI using our Universal Encoder before being decoded. Both successful reconstructions and failure cases are presented. As observed, the reconstructions from the encoded fMRI are on par with those from the GT fMRI, despite the decoder being trained solely on GT fMRI. This highlights the effectiveness of our Universal Encoder, demonstrating that the encoded fMRI retains information exists in the GT fMRI.

### v3

@@ A.2 Training and Lora Modification
The Lora adaptation  (hu2021lora) is done by adding learned low rank matrices to weights in the original network. In the original paper the value projection and query projection weights ($W_{v}$$W_{q}$) in the self attention block are modified. We only modify the output projection weights($W_{o}$) of the self attention block.
@@ Appendix C Statistical Tests
Additionally, we tested the prediction correlations of individual voxels (across all subjects) against a null hypothesis. The null hypothesis was based on the distribution of correlations between pairs of independent Gaussian random vectors of size N = 982, matching the number of image-fMRI pairs in the test set. The one-sided statistical significance was estimated by comparing the predicted voxel correlations for each subject on the test set with those from the null distribution. We applied a statistical threshold of P << 0.05, correcting for multiple comparisons using the FDR procedure. Our results indicate that the majority of voxel predictions by the Universal Encoder are statistically significant, with an average of 98% across all subjects. Detailed results in Table S1.
@@ Appendix D Additional Figures
Figure S8: Comparison against additional Encoder models. Pearson correlation between predicted and ground-truth fMRI of Subject1 from NSD dataset. The figure presents comparisons across three prominent encoders from different papers (gaziv2022self; wang2023better; takagi2023high), as well as the universal encoder trained on both single-subject and multi-subject data. (a) Shows the median values along with the 25th and 75th percentiles across the five models. (b) Compares correlation scores across different coarse brain regions (ROIs) for the same five models. .
@@
Figure S11: Retrieval Results: Our universal encoder trained on multiple subjects significantly improves retrieval accuracy over both subject-specific models as evidenced by the calculated p-value. We conducted 10 retrieval experiments, wherein we randomly sampled 999 different distracted images for each “query” fMRI. Results are compared across three models: (i) ”Baseline Encoder” – the encoder of (gaziv2022self) trained separately on each subject, (ii) ”Universal Encoder - Single Subject” – our architecture trained separately on each subject, and (iii) ”Universal Encoder - Multiple Subjects” – our model trained on combined data from 8 subjects. In (a) we present the mean top-1 accuracy (across the 10 experiments) along with the standard deviation for each model and subject. (b) depicts the mean top-5 accuracy along with the standard deviation for each model and subject.
Figure S12: MSE Evaluation - The Wisdom of a Crowd of Brains. By aggregating data from multiple subjects, our Universal-Encoder improves the encoding of any individual subject. We present results calculating the MSE between predicted and ground-truth fMRI (median value with 25th and 75th percentiles). We compared 3 models: (i) The ”Baseline” single-subject encoder of (gaziv2022self), (ii) ”Universal Encoder - single subject” – our architecture trained on each subject separately, (iii) ”Universal Encoder - multiple subjects” – our model trained on data from 8 subjects.
@@
Figure S13: Qualitative Evaluation of Image Reconstruction from fMRI. Visual comparison of image reconstruction from fMRI for Subject 1 from the NSD dataset. The figure showcases the ground truth (GT) image seen by the subject, alongside reconstructions generated using the fMRI-to-Image decoding model (MindEye (scotti2024reconstructing)) trained on Subject 1’s fMRI data. Both the decoder and our Universal Encoder were trained on the same dataset and evaluated on the same test set. Reconstructions were evaluated on the test set using two inputs: (1) ”Reconstruction from GT fMRI” where the decoder processes the real fMRI recordings of the test images, and (2) ”Reconstruction from our encoded fMRI”, where the test images are first encoded into predicted fMRI using our Universal Encoder before being decoded. Both successful reconstructions and failure cases are presented. As observed, the reconstructions from the encoded fMRI are on par with those from the GT fMRI, despite the decoder being trained solely on GT fMRI. This highlights the effectiveness of our Universal Encoder, demonstrating that the encoded fMRI retains information exists in the GT fMRI.